# On the singular loci of higher secant varieties of Veronese embeddings

### v2

@@ 5.2. Singularity and non-singularity
and $\operatorname{rank}\phi(f_{5})=12$. The conormal space $\hat{N}^{\ast}_{f_{5}}\sigma_{4}(v_{3}({\mathbb{P}}^{3}))$ contains a 4-dimensional subspace corresponding to $\langle-y_{0}y_{3}^{2}+4y_{1}y_{2}y_{3}-24y_{2}^{3},-y_{1}y_{3}^{2}+12y_{2}^{2}y_{3},y_{2}y_{3}^{2},y_{3}^{3}\rangle$ which can be spanned by $\{y_{3}\otimes x_{0},2y_{1}\otimes x_{0}+12y_{2}\otimes x_{1}+y_{3}\otimes x_{2}\}$ in $\ker\phi(f_{5})\subset U^{\ast}\otimes U$ and
@@
$+(d-1)x_{0}^{d-2}\otimes x_{0}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{3}\big{)}+(d-1)x_{0}^{d-2}\otimes x_{3}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{0}\big{)}~{}$
@@
Then, one can check that $A_{2}\cdot B_{2}$ produces a subspace of $\hat{N}^{\ast}_{f_{5}}\sigma_{4}(v_{3}({\mathbb{P}}^{3}))$ in ${S}^{d}U^{\ast}$ which is the degree $d$-part of an ideal $I_{2}$ generated by 19 quartics
$\{~{}36y_{0}^{2}y_{2}^{2}-12(d-2)y_{0}y_{1}^{2}y_{2}+(d-2)^{2}y_{1}^{4},~{}6y_{0}^{2}y_{2}y_{3}-(d-2)y_{0}y_{1}^{2}y_{3}-6(d-1)y_{0}y_{1}y_{2}^{2}+(d-1)(d-2)y_{1}^{3}y_{2},$
$~{}-y_{0}^{2}y_{3}^{2}+2(d-1)y_{0}y_{1}y_{2}y_{3}-(d-1)^{2}y_{1}^{2}y_{2}^{2},~{}-6y_{0}y_{1}y_{2}y_{3}+(d-2)y_{1}^{3}y_{3},~{}-y_{0}y_{1}y_{3}^{2}+(d-1)y_{1}^{2}y_{2}y_{3},$
$~{}-6y_{0}y_{2}^{3}+(d-2)y_{1}^{2}y_{2}^{2},~{}y_{0}y_{2}^{2}y_{3}-(d-1)y_{1}y_{2}^{3},~{}y_{0}y_{2}y_{3}^{2}-(d-1)y_{1}y_{2}^{2}y_{3},~{}y_{0}y_{3}^{3}-(d-1)y_{1}y_{2}y_{3}^{2},$
$~{}(d-2)y_{1}^{2}y_{2}y_{3}-6(d-1)y_{1}y_{2}^{3},~{}y_{1}^{2}y_{3}^{2},~{}y_{1}y_{2}^{2}y_{3},~{}y_{1}y_{2}y_{3}^{2},~{}y_{1}y_{3}^{3},~{}y_{2}^{4},~{}y_{2}^{3}y_{3},y_{2}^{2}y_{3}^{2},~{}y_{2}y_{3}^{3},~{}y_{3}^{4}\}~{}.$
@@
As a direct consequence of main results in the paper, we also obtain the following corollary on (non-)singularity of subsecant loci in the 4-th secant variety.
@@
    $\sigma_{4}(v_{d}({\mathbb{P}}^{1}))\setminus\sigma_{3}(v_{d}({\mathbb{P}}^{n}))\subset\operatorname{Sm}(\sigma_{4}(v_{d}({\mathbb{P}}^{n})))$ if $d\geq 7$. $\sigma_{4}(v_{d}({\mathbb{P}}^{1}))$ is non-trivial singular locus when $d=6$ and $\sigma_{4}(v_{d}({\mathbb{P}}^{1}))\subset\sigma_{3}(v_{d}({\mathbb{P}}^{n}))$ when $d\leq 5$.
@@
(i) For $m=2$, Theorem 2 (ii) tells $\sigma_{4}(v_{d}({\mathbb{P}}^{m}))$ is non-trivial singular locus in $\sigma_{4}(v_{d}({\mathbb{P}}^{n}))$ if $d=3,n\geq 4$. The case $(d,n)=(3,3)$ is also discussed in Theorem 2 (iv).
When $d=4$, we can say that a general point in $\sigma_{4}(v_{d}({\mathbb{P}}^{2}))\setminus\sigma_{3}(v_{d}({\mathbb{P}}^{n}))$ is smooth by Theorem 3 (i). For any $d\geq 5$, the same conclusion comes from Theorem 2 (i).
@@
Let $V$ be a $(n+1)$-dimensional complex vector space $(n\geq 1)$ and $v_{d}({\mathbb{P}}V)\subset{\mathbb{P}}^{N}$ be the image of the $d$-uple $(d\geq 2)$ Veronese embedding of ${\mathbb{P}}V$. Assume that $(k,d,n)$ satisfies one of the following conditions:
@@
(29) $\big{\{}f\in{\mathbb{P}}S^{d}V~{}|~{}\textrm{$f$ is any form which can be expressed using at most $k-1$ variables}\big{\}}~{},$
which is an irreducible locus of dimension $(k-1)(n-k+2)+{d+k-2\choose d}-1$ and is equal to the maximum subsecant locus $\Sigma_{k,d}(\min\{k-1,n\}-1;{\mathbb{P}}V)$.
@@
For case (i), the assertion is immediate since it corresponds to symmetric matrices. In case (ii), we draw the conclusion from the fact $\operatorname{Sing}(\sigma_{2}(v_{d}({\mathbb{P}}V)))=v_{d}({\mathbb{P}}V)$ for every $d,n$ [20].
@@
(30) $\sigma_{k-1}(v_{d}({\mathbb{P}}V))\subset\bigcup_{{\mathbb{P}}^{k-2}\subset{\mathbb{P}}V}\langle\,v_{d}({\mathbb{P}}^{k-2})\,\rangle~{}.$
We note that the right hand side of (30) is an irreducible and closed subvariety of $\sigma_{k}(v_{d}({\mathbb{P}}V))$, since it coincides with a subvariety $\bigcup_{\Lambda\in\mathrm{Im}\Phi}\Lambda$ where a map $\Phi:\mathbb{G}(k-2,n)\to\mathbb{G}({d+k-2\choose d}-1,N)$ sending each subspace $L$ of dimension $k-2$ to the linear span $\langle\,v_{d}(L)\,\rangle$ in ${\mathbb{P}}^{N}$ is regular (see e.g. [17, example 6.10, proposition 6.13]). Then, because a general element of the left hand side is of the form $\ell_{1}^{d}+\cdots+\ell_{k-1}^{d}$ for some linear forms $\ell_{i}$’s, it belongs to $\langle\,v_{d}({\mathbb{P}}^{k-2})\,\rangle$ for some ${\mathbb{P}}^{k-2}\subset{\mathbb{P}}V$ so that the closure is also contained in the subvariety $\bigcup_{{\mathbb{P}}^{k-2}\subset{\mathbb{P}}V}\langle\,v_{d}({\mathbb{P}}^{k-2})\,\rangle$.
For case (iii), by [16, theorem 2.1, remark 2.4(a), and corollary 2.11] and Theorem 1, and for case (iv), by Theorem 5, we know that
$\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}V)))=\sigma_{k-1}(v_{d}({\mathbb{P}}V))\cup\big{\{}\bigcup_{{\mathbb{P}}^{k-2}\subset{\mathbb{P}}V}\sigma_{k}(v_{d}({\mathbb{P}}^{k-2}))\big{\}}~{},$
which can be also written as $\sigma_{k-1}(v_{d}({\mathbb{P}}V))\cup\Sigma_{k,d}(\min\{k-1,n\}-1;{\mathbb{P}}V)$. In both case (iii) and (iv), we have $\sigma_{k}(v_{d}({\mathbb{P}}^{k-2}))=\langle\,v_{d}({\mathbb{P}}^{k-2})\,\rangle$. Thus, by the claim above, the singular locus is equal to $\bigcup_{{\mathbb{P}}^{k-2}\subset{\mathbb{P}}V}\langle\,v_{d}({\mathbb{P}}^{k-2})\,\rangle=\Sigma_{k,d}(\min\{k-1,n\}-1;{\mathbb{P}}V)$, which is irreducible and can be described as written in the statement. The formula for dimension is immediate from dimension counting. ∎
@@ 6. Concluding remark
So far, we reported results on singular loci of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ coming from the subsecant loci. To the best of our knowledge, there has been no general idea or clear consensus on the singular locus of arbitrary higher secant variety of any Veronese variety yet. From this point of view, the present paper contributes to it as making a more visible picture on the singular locus via showing a generic smoothness of the subsecant loci for relatively low $k$’s and confirming singularity of the same loci for other $k$’s.
As we mentioned in the introduction, each point $p\in\sigma_{k}(v_{d}({\mathbb{P}}^{n}))\setminus\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ is located in $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\setminus\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ for some $1\leq m\leq\min\{k-1,n\}$. To make the picture more complete, we have two future issues: (i) On the subsecant loci (i.e. $m<\min\{k-1,n\}$), one needs to check (non-)singularity not only at a general point but also at every point and (ii) Points in the full-secant locus (i.e. $m=\min\{k-1,n\}$) should be treated.
Issue (i) is expected to be very complicated, because at some special point a singularity can also happen even for a low $k$ as shown in Example 27 (in fact, we can generate more examples using a similar idea). For the points in the subsecant loci, in general, one could not hope to find some nice ‘normal forms’ and the situation is expected to be wild (in other words, the subsecant loci may not be covered with finitely many nice families of $\mathbf{SL}$-orbits). But, still we can push on our viewpoint a bit further and, along the same spirit, we can refine a main result of this paper in the following manner. Based on the singularity results in Theorems 2 and 3 and using a similar estimation as §2.3, more generally we have:
@@
Suppose that $k>\mu$ in case of Theorem 2 or $k\geq\lfloor\mu\rfloor$ in case of Theorem 3 (in other words, the subsecant variety $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ is a singular locus in $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$). Let $r\leq n-m$. Then, the following does hold
(31) $J(\sigma_{k}(v_{d}({\mathbb{P}}^{m})),\sigma_{r}(v_{d}({\mathbb{P}}^{n})))\subset\operatorname{Sing}(\sigma_{k+r}(v_{d}({\mathbb{P}}^{n})))~{},$
where $J(X,Y)$ denotes the ‘(embedded) join’ of two subvarieties $X,Y$ in some ${\mathbb{P}}^{N}$.
@@
Suppose that the inclusion (31) does not hold. Then, taking $x_{1},\dots,x_{k}$ to be general points of $v_{d}({\mathbb{P}}^{m})$ and $x_{k+1},\dots,x_{k+r}$ to be general points of $v_{d}({\mathbb{P}}^{n})$, we may assume that $x\notin\operatorname{Sing}(\sigma_{k+r}(v_{d}({\mathbb{P}}^{n})))$ for a general $x\in\langle\,x_{1},\dots,x_{k+r}\,\rangle$. By Terracini lemma, we have
(32) $L_{1}:=\langle\,\mathbb{T}_{x_{1}}v_{d}({\mathbb{P}}^{n}),\dots,\mathbb{T}_{x_{k}}v_{d}({\mathbb{P}}^{n})\,\rangle\subset\mathbb{T}_{x}\sigma_{k+r}(v_{d}({\mathbb{P}}^{n}))$
and by the assumption on $k$ we know $\dim L_{1}>kn+k-1$.
@@
Moreover, we may assume $L_{1}\cap L_{2}=\emptyset$ as follows. Taking ${\mathbb{P}}^{n-m-1}\subset{\mathbb{P}}^{n}$ such that $x_{k+1},\dots,x_{k+r}\in{\mathbb{P}}^{r-1}\subset{\mathbb{P}}^{n-m-1}$ and ${\mathbb{P}}^{m}\cap{\mathbb{P}}^{n-m-1}=\emptyset$, and changing coordinates $t_{0},\dots,t_{m},u_{1},\dots,u_{m^{\prime}}$ on ${\mathbb{P}}^{n}$ as in §2.3, we may say that ${\mathbb{P}}^{n-m-1}$ is the zero set of $t_{0}=\dots=t_{m}=0$ and ${\mathbb{P}}^{m}$ is the zero set of $u_{1}=\dots=u_{m^{\prime}}=0$. For a point $x^{\prime}\in{\mathbb{P}}^{m}$, using the parametrization (11), the tangent space $\mathbb{T}_{v_{d}(x^{\prime})}v_{d}({\mathbb{P}}^{n})$ is spanned by the rows of the matrix of the form $[\ast:\mathsf{O}]$ as (13). On the other hand, for a point $x^{\prime\prime}\in{\mathbb{P}}^{n-m-1}$ and for an affine open set containing $x^{\prime\prime}$ we may take $u_{m^{\prime}}=1$ instead of $t_{0}=1$. Then, the only part in the parametrization of $v_{d}$ which contributes $\mathbb{T}_{v_{d}(x^{\prime\prime})}v_{d}({\mathbb{P}}^{n})$ is
@@
This new singular locus $J(\sigma_{k}(v_{d}({\mathbb{P}}^{m})),\sigma_{r}(v_{d}({\mathbb{P}}^{n})))$ in (31) can be seen as a ‘partial version’ of subsecant locus in this paper. In particular, it contains $m$-subsecant variety $\sigma_{k+r}(v_{d}({\mathbb{P}}^{m}))=J(\sigma_{k}(v_{d}({\mathbb{P}}^{m})),\sigma_{r}(v_{d}({\mathbb{P}}^{m})))$. So, let us call such a locus a partial subsecant locus of $\sigma_{k+r}(v_{d}({\mathbb{P}}^{n}))$. We note that the singularity of a specific form $f=x^{2}y^{2}+z^{4}$ in Example 27 can be explained using this notion; $f$ is a point of $\Sigma_{4,4}(2;{\mathbb{P}}^{3})$ where only a generic smoothness is known by Theorems 3 (i), but $f$ also belongs to a partial subsecant locus $J(\sigma_{3}(v_{4}({\mathbb{P}}^{1})),\sigma_{1}(v_{4}({\mathbb{P}}^{3})))$ which is singular by Theorem 33.
@@
Note that the answer to Question 35 is affirmative in cases of $k=2$ classically and $k=3$ (by [16]) and $k=4$ (by Theorem 29). For a large value $k$ compared to $n$ (e.g. $n<k-1$), Question 35 may be answered negatively as in the following example.
@@
Let us consider $\sigma_{14}(v_{8}({\mathbb{P}}^{2}))$, the $14$-th secant variety of the Veronese variety $v_{8}({\mathbb{P}}^{2})$. Take general 14 points on $v_{8}({\mathbb{P}}^{2})$. In [3, remark 4.10], the authors presented a concrete point in the linear span of the 14 points which is a non-normal point to $\sigma_{14}(v_{8}({\mathbb{P}}^{2}))$. Note that one can also check this singular point does not belong to $\mathcal{D}$, the locus of all partial subsecants.
@@
“ For $\vec{\mathbf{n}}:=(n_{1},n_{2},\ldots,n_{r})$, let $X$ be the Segre embedding ${\mathbb{P}}^{n_{1}}\times{\mathbb{P}}^{n_{2}}\times\cdots\times{\mathbb{P}}^{n_{r}}\subset{\mathbb{P}}^{\prod(n_{i}+1)-1}=:{\mathbb{P}}^{\beta(\vec{\mathbf{n}})}$ and denote $\sigma_{k}(X)$ by $\sigma_{k}(\vec{\mathbf{n}})$, $\dim\sigma_{k}(X)$ by $s_{k}(\vec{\mathbf{n}})$. For every $\vec{\mathbf{m}}=(m_{1},m_{2},\ldots,m_{r})$ with $0\leq m_{i}\leq\min\{k-1,n_{i}-1\}$, besides a few exceptional cases, we have that the following holds
@@
which can recover the result on the singular locus of the secant varieties of Segre embeddings ([27, corollary 7.17]) for $k=2$. Note that, if we assume that everything is non-defective, then the ranges above can be computed as

### v3

@@ 5.2. Singularity and non-singularity
and $\operatorname{rank}\phi(f_{5})=12$. The conormal space $\hat{N}^{\ast}_{[f_{5}]}\sigma_{4}(v_{3}({\mathbb{P}}^{3}))$ contains a 4-dimensional subspace corresponding to $\langle-y_{0}y_{3}^{2}+4y_{1}y_{2}y_{3}-24y_{2}^{3},-y_{1}y_{3}^{2}+12y_{2}^{2}y_{3},y_{2}y_{3}^{2},y_{3}^{3}\rangle$ which can be spanned by $\{y_{3}\otimes x_{0},2y_{1}\otimes x_{0}+12y_{2}\otimes x_{1}+y_{3}\otimes x_{2}\}$ in $\ker\phi(f_{5})\subset U^{\ast}\otimes U$ and
@@
$+(d-1)x_{0}^{d-2}\otimes x_{0}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{3}\big{)}+(d-1)x_{0}^{d-2}\otimes x_{3}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{0}\big{)}$
@@
Then, one can check that $A_{2}\cdot B_{2}$ produces a subspace of $\hat{N}^{\ast}_{[f_{5}]}\sigma_{4}(v_{3}({\mathbb{P}}^{3}))$ in ${S}^{d}U^{\ast}$ which is the degree $d$-part of an ideal $I_{2}$ generated by 19 quartics
[…]
$-y_{0}^{2}y_{3}^{2}+2(d-1)y_{0}y_{1}y_{2}y_{3}-(d-1)^{2}y_{1}^{2}y_{2}^{2},~{}-6y_{0}y_{1}y_{2}y_{3}+(d-2)y_{1}^{3}y_{3},~{}-y_{0}y_{1}y_{3}^{2}+(d-1)y_{1}^{2}y_{2}y_{3},$
$-6y_{0}y_{2}^{3}+(d-2)y_{1}^{2}y_{2}^{2},~{}y_{0}y_{2}^{2}y_{3}-(d-1)y_{1}y_{2}^{3},~{}y_{0}y_{2}y_{3}^{2}-(d-1)y_{1}y_{2}^{2}y_{3},~{}y_{0}y_{3}^{3}-(d-1)y_{1}y_{2}y_{3}^{2},$
$\,(d-2)y_{1}^{2}y_{2}y_{3}-6(d-1)y_{1}y_{2}^{3},~{}y_{1}^{2}y_{3}^{2},~{}y_{1}y_{2}^{2}y_{3},~{}y_{1}y_{2}y_{3}^{2},~{}y_{1}y_{3}^{3},~{}y_{2}^{4},~{}y_{2}^{3}y_{3},y_{2}^{2}y_{3}^{2},~{}y_{2}y_{3}^{3},~{}y_{3}^{4}\ \}.$
@@
As a direct consequence of the main results in the paper, we also obtain the following corollary on the (non-)singularity of subsecant loci in the 4-th secant variety.
@@
    $\sigma_{4}(v_{d}({\mathbb{P}}^{n}))$ is smooth at each point in $\sigma_{4}(v_{d}({\mathbb{P}}^{1}))\setminus\sigma_{3}(v_{d}({\mathbb{P}}^{n}))$ if $d\geq 7$. $\sigma_{4}(v_{d}({\mathbb{P}}^{1}))$ is non-trivial singular locus when $d=6$ and $\sigma_{4}(v_{d}({\mathbb{P}}^{1}))\subset\sigma_{3}(v_{d}({\mathbb{P}}^{n}))$ in case of $d\leq 5$.
@@
(i) For $m=2$, Theorem 2 (ii) says that $\sigma_{4}(v_{d}({\mathbb{P}}^{m}))$ is a non-trivial singular locus in $\sigma_{4}(v_{d}({\mathbb{P}}^{n}))$ if $d=3,n\geq 4$. The case $(d,n)=(3,3)$ is also discussed in Theorem 2 (iv).
When $d=4,5$ and $6$, we can say that a general point in $\sigma_{4}(v_{d}({\mathbb{P}}^{2}))\setminus\sigma_{3}(v_{d}({\mathbb{P}}^{n}))$ is smooth by Theorem 3 (i). For any $d\geq 7$, the same conclusion follows from Theorem 2 (i).
@@
Let $V$ be an $(n+1)$-dimensional complex vector space $(n\geq 1)$ and $v_{d}({\mathbb{P}}V)\subset{\mathbb{P}}^{N}$ be the image of the $d$-uple $(d\geq 2)$ Veronese embedding of ${\mathbb{P}}V$. Assume that $(k,d,n)$ satisfies one of the following conditions:
@@
(37) $\big{\{}\,f\in{\mathbb{P}}S^{d}V~{}|~{}\textrm{$f$ is any form which can be expressed using at most $k-1$ variables}\,\big{\}},$
which is an irreducible locus of dimension
$(k-1)(n-k+2)+{d+k-2\choose d}-1$
and is equal to the maximum subsecant locus $\Sigma_{k,d}(\min\{k-1,n\}-1;{\mathbb{P}}V)$.
@@
For case (i), the assertion is immediate since it corresponds to symmetric matrices. In case (ii), we draw the conclusion from the fact that $\operatorname{Sing}(\sigma_{2}(v_{d}({\mathbb{P}}V)))=v_{d}({\mathbb{P}}V)$ for every $d,n$ [Kan99].
@@
(38) $\sigma_{k-1}(v_{d}({\mathbb{P}}V))\subset\bigcup_{{\mathbb{P}}^{k-2}\subset{\mathbb{P}}V}\langle\,v_{d}({\mathbb{P}}^{k-2})\,\rangle.$
We note that the right hand side of (38) is an irreducible and closed subvariety of $\sigma_{k}(v_{d}({\mathbb{P}}V))$, since it coincides with a subvariety $\bigcup_{\Lambda\in\mathrm{Im}\Phi}\Lambda$, where a map
$\Phi:\mathbb{G}(k-2,n)\to\mathbb{G}\left(\binom{d+k-2}{d}-1,N\right)$
sending each subspace $L$ of dimension $k-2$ to the linear span $\langle\,v_{d}(L)\,\rangle$ in ${\mathbb{P}}^{N}$ is regular (see e.g. [Harr92, Example 6.10, Proposition 6.13]). Then, because a general element of the left hand side is of the form $\ell_{1}^{d}+\cdots+\ell_{k-1}^{d}$ for some linear forms $\ell_{i}$’s, it belongs to $\langle\,v_{d}({\mathbb{P}}^{k-2})\,\rangle$ for some ${\mathbb{P}}^{k-2}\subset{\mathbb{P}}V$ so that the closure is also contained in the subvariety $\bigcup_{{\mathbb{P}}^{k-2}\subset{\mathbb{P}}V}\langle\,v_{d}({\mathbb{P}}^{k-2})\,\rangle$.
For case (iii), by [Han18, Theorem 2.1, Remark 2.4(a), and Corollary 2.11] and Theorem 1, and for case (iv), by Theorem 5, we know that
$\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}V)))=\sigma_{k-1}(v_{d}({\mathbb{P}}V))\cup\big{\{}\bigcup_{{\mathbb{P}}^{k-2}\subset{\mathbb{P}}V}\sigma_{k}(v_{d}({\mathbb{P}}^{k-2}))\big{\}},$
which can also be written as
$\sigma_{k-1}(v_{d}({\mathbb{P}}V))\cup\Sigma_{k,d}(\min\{k-1,n\}-1;{\mathbb{P}}V).$
In both cases (iii) and (iv), we have $\sigma_{k}(v_{d}({\mathbb{P}}^{k-2}))=\langle\,v_{d}({\mathbb{P}}^{k-2})\,\rangle$. Thus, by the above claim, the singular locus is equal to
$\bigcup_{{\mathbb{P}}^{k-2}\subset{\mathbb{P}}V}\langle\,v_{d}({\mathbb{P}}^{k-2})\,\rangle=\Sigma_{k,d}(\min\{k-1,n\}-1;{\mathbb{P}}V),$
which is irreducible and can be described as written in the statement. The formula for the dimension is immediate from dimension counting. ∎
@@ 6. Concluding remark
So far, we have reported results on singular loci of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ coming from the subsecant loci. To the best of our knowledge, there is no general idea or clear consensus on the singular locus of an arbitrary higher secant variety of any Veronese variety yet. From this point of view, the present paper contributes by providing a more visible picture on the singular locus via showing a generic smoothness of the subsecant loci for relatively low $k$’s and confirming the singularity of the same loci for other $k$’s.
As we mentioned in the introduction, each point $p\in\sigma_{k}(v_{d}({\mathbb{P}}^{n}))\setminus\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ is located in $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\setminus\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ for some $1\leq m\leq\min\{k-1,n\}$. To make the picture more complete, we have two future issues: (i) On the subsecant loci (i.e., $m<\min\{k-1,n\}$), one needs to check the (non-)singularity not only at a general point but also at every point and (ii) Points in the full-secant locus (i.e., $m=\min\{k-1,n\}$) should be treated.
Issue (i) is expected to be very complicated, because at some special point a singularity can also occur even for a low $k$ as shown in Example 27 (in fact, we can generate more examples using a similar idea). For the points in the subsecant loci, in general, one could not hope to find some nice ‘normal forms’ and the situation is expected to be wild (in other words, the subsecant loci may not be covered with finitely many nice families of $\mathbf{SL}$-orbits). But, still we can push on our viewpoint a bit further and, along the same spirit, we can refine a main result of this paper in the following manner. Based on the singularity results in Theorems 1, 2, and 3 and using a similar estimation as in §2.3, more generally we have:
@@
Suppose that $m=1$ and $k,d$ satisfy the condition of (ii) or (iii) of Theorem 1, or suppose that $k,d,m$ satisfy the condition of (ii) or (iii) of Theorem 2 or Theorem 3; in other words, the $m$-subsecant variety $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ is a singular locus in $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$. Let $11\leq m\leq n-$ 1 and $r\leq n-m$. Then, unless $\sigma_{k+r}(v_{d}(\mathbb{P}^{n}))$ fills up the ambient space $\mathbb{P}^{N}$, the following holds
(39) $J(\sigma_{k}(v_{d}({\mathbb{P}}^{m})),\sigma_{r}(v_{d}({\mathbb{P}}^{n})))\subset\operatorname{Sing}(\sigma_{k+r}(v_{d}({\mathbb{P}}^{n}))),$
where $J(X,Y)$ denotes the ‘(embedded) join’ of two subvarieties $X,Y$ in their ambient space.
@@
Suppose that the inclusion (39) does not hold. Then, taking $x_{1},\dots,x_{k}$ to be general points of $v_{d}({\mathbb{P}}^{m})$ and $x_{k+1},\dots,x_{k+r}$ to be general points of $v_{d}({\mathbb{P}}^{n})$, we may assume that $x\notin\operatorname{Sing}(\sigma_{k+r}(v_{d}({\mathbb{P}}^{n})))$ for a general $x\in\langle\,x_{1},\dots,x_{k+r}\,\rangle$. By Terracini’s lemma, we have
(40) $L_{1}=\langle\,\mathbb{T}_{x_{1}}v_{d}({\mathbb{P}}^{n}),\dots,\mathbb{T}_{x_{k}}v_{d}({\mathbb{P}}^{n})\,\rangle\subset\mathbb{T}_{x}\sigma_{k+r}(v_{d}({\mathbb{P}}^{n}))$
and by the assumption on $k$ we know that $\dim L_{1}>kn+k-1$.
@@
Moreover, we may assume $L_{1}\cap L_{2}=\emptyset$ as follows. Taking ${\mathbb{P}}^{n-m-1}\subset{\mathbb{P}}^{n}$ such that $x_{k+1},\dots,x_{k+r}\in{\mathbb{P}}^{r-1}\subset{\mathbb{P}}^{n-m-1}$ and ${\mathbb{P}}^{m}\cap{\mathbb{P}}^{n-m-1}=\emptyset$, and changing coordinates $t_{0},\dots,t_{m},u_{1},\dots,u_{m^{\prime}}$ on ${\mathbb{P}}^{n}$ as in §2.3, we may say that ${\mathbb{P}}^{n-m-1}$ is the zero set of $t_{0}=\dots=t_{m}=0$ and ${\mathbb{P}}^{m}$ is the zero set of $u_{1}=\dots=u_{m^{\prime}}=0$. For a point $x^{\prime}\in{\mathbb{P}}^{m}$, using the parameterization (16), the tangent space $\mathbb{T}_{v_{d}(x^{\prime})}v_{d}({\mathbb{P}}^{n})$ is spanned by the rows of the matrix of the form $[\ast:\mathsf{O}]$ as (18). On the other hand, for a point $x^{\prime\prime}\in{\mathbb{P}}^{n-m-1}$ and for an affine open set containing $x^{\prime\prime}$ we may take $u_{m^{\prime}}=1$ instead of $t_{0}=1$. Then, the only part on the parameterization of $v_{d}$ which contributes $\mathbb{T}_{v_{d}(x^{\prime\prime})}v_{d}({\mathbb{P}}^{n})$ is
@@
This new singular locus $J(\sigma_{k}(v_{d}({\mathbb{P}}^{m})),\sigma_{r}(v_{d}({\mathbb{P}}^{n})))$ in (39) can be seen as a ‘partial version’ of subsecant locus in this paper. In particular, it contains the $m$-subsecant variety $\sigma_{k+r}(v_{d}({\mathbb{P}}^{m}))=J(\sigma_{k}(v_{d}({\mathbb{P}}^{m})),\sigma_{r}(v_{d}({\mathbb{P}}^{m})))$. So, let us call such a locus a partial subsecant locus of $\sigma_{k+r}(v_{d}({\mathbb{P}}^{n}))$. We note that the singularity of a specific form $f=x^{2}y^{2}+z^{4}$ in Example 27 can be explained using this notion; $f$ is a point of $\Sigma_{4,4}(2;{\mathbb{P}}^{3})$ where only a generic smoothness is known by Theorems 3 (i), but $f$ also belongs to a partial subsecant locus $J(\sigma_{3}(v_{4}({\mathbb{P}}^{1})),\sigma_{1}(v_{4}({\mathbb{P}}^{3})))$ which is singular by Theorem 33.
@@
Note that the answer to Question 35 is affirmative in cases of $k=2$ classically and $k=3$ (by [Han18]) and $k=4$ (by Theorem 29). For a large value $k$ compared to $n$ (e.g. $n<k-1$), Question 35 may be answered negatively as in the following example.
@@
Let us consider $\sigma_{14}(v_{8}({\mathbb{P}}^{2}))$, the $14$-th secant variety of the Veronese variety $v_{8}({\mathbb{P}}^{2})$. Take general 14 points on $v_{8}({\mathbb{P}}^{2})$. In [AC20, remark 4.10], the authors presented a concrete point in the linear span of the 14 points which is a non-normal point to $\sigma_{14}(v_{8}({\mathbb{P}}^{2}))$. Note that one can also check this singular point does not belong to $\mathcal{D}$, the locus of all partial subsecants.
@@
“ For $\vec{\mathbf{n}}=(n_{1},n_{2},\ldots,n_{r})$, let $X$ be the Segre embedding ${\mathbb{P}}^{n_{1}}\times{\mathbb{P}}^{n_{2}}\times\cdots\times{\mathbb{P}}^{n_{r}}\subset{\mathbb{P}}^{\prod(n_{i}+1)-1}=:{\mathbb{P}}^{\beta(\vec{\mathbf{n}})}$ and denote $\sigma_{k}(X)$ by $\sigma_{k}(\vec{\mathbf{n}})$, the expected dimension of $\sigma_{k}(\vec{\mathbf{n}})$ by $s_{k}(\vec{\mathbf{n}})$. For every $\vec{\mathbf{m}}=(m_{1},m_{2},\ldots,m_{r})$ with $0\leq m_{i}\leq\min\{k-1,n_{i}-1\}$, besides a few exceptional cases, we have that the following holds
@@
which can recover the result on the singular locus of the secant varieties of Segre embeddings ([MOZ15, Corollary 7.17]) for $k=2$. Note that, if we assume that everything is non-defective, then the ranges above can be computed as